Electronic Components Defect Probability Analysis

00:00 8 marks

  • Build Confidence

A factory produces electronic components. It is known that 3% of components produced by machine A are defective and 4% of components produced by machine B are defective. Machine A produces 60% of the total components. A component is selected at random.

(a) Find the probability that the selected component is defective. [3]
(b) Given that the selected component is defective, find the probability that it was produced by machine B. [3]
(c) It is later found that 1% of defective components are incorrectly classified as non-defective by the quality control test, and 2% of non-defective components are incorrectly classified as defective. Find the probability that a component which is classified as defective is actually non-defective. [2]

Part a

3 minutes 3 marks

Question

Find the probability that the selected component is defective. [3]



Hint Hint

Hint

Use the law of total probability with P(D|A)=0.03, P(D|B)=0.04, and P(A)=0.6 (hence P(B)=0.4).

  • Answer Answer
  • Mark Scheme Mark Scheme
  • Explanation Explanation
  • Answer Answer

Complete Answer

P(D) = 0.034

  • Mark Scheme Mark Scheme
  • Answer Explanation Answer Explanation

Part b

3 minutes 3 marks

Question

Given that the selected component is defective, find the probability that it was produced by machine B. [3]



Hint Hint

Hint

Apply Bayes’ theorem using P(D|B), P(B), and the result from part (a).

  • Answer Answer
  • Mark Scheme Mark Scheme
  • Explanation Explanation
  • Answer Answer

Complete Answer

P(B|D) ≈ 0.471

  • Mark Scheme Mark Scheme
  • Answer Explanation Answer Explanation

Part c

4 minutes 2 marks

Question

It is later found that 1% of defective components are incorrectly classified as non-defective, and 2% of non-defective components are incorrectly classified as defective. Find the probability that a component which is classified as defective is actually non-defective. [2]



Hint Hint

Hint

Consider P(classified defective|defective)=0.99, P(classified defective|non-defective)=0.02, and note that P(non-defective)=1-P(D)=0.966.

  • Answer Answer
  • Mark Scheme Mark Scheme
  • Explanation Explanation
  • Answer Answer

Complete Answer

Approximately 0.365

  • Mark Scheme Mark Scheme
  • Answer Explanation Answer Explanation

Math Editor

Report Question


Please enter keywords